OpenStudy (anonymous):

Josie makes fruit punch by mixing fruit juice and lemonade in ratio of 2:7 She needs to make 63 litres of punch for a party. How much of each ingredient does she need? Fruit Juice: 14 litres Lemonade: 49 litres (tell me if wrong) She wants more though, She has 10 litres of fruit juice left and plenty of lemonade. How much extra punch can she make? im confused on how to answer the second one so can you explain how you got the answer

OpenStudy (anonymous):

get*

OpenStudy (anonymous):

if she's got 10 liters of juice and plenty of lemonade,then she uses (10*7/2)=35 liters of lemonade and ends up with 45 liters of punch..

OpenStudy (anonymous):

why do you divide it by 2?

OpenStudy (anonymous):

i just took the ratio:(juice/lemonade)=(2/7),juice =10liters...

OpenStudy (anonymous):

ok thanks